OpenStudy (anonymous):

Factor 9x^6 – 16y^6 completely. First factor each term. 9x^6 = (__)^2 16y^6 = (__)^2

OpenStudy (cggurumanjunath):

(3x^3)^2=9x^6

OpenStudy (cggurumanjunath):

(4y^3)^2=16y^6

OpenStudy (anonymous):

Ok thanks. Then use a^2 – b^2 = (a – b)(a + b). Show your work. 9x^6 – 16y^6 =

OpenStudy (anonymous):

would it be 3 - 4 and 3 + 4 @ CGGURUMANJUNATH ?

OpenStudy (cggurumanjunath):

(3x^3-4y^^3)^2

OpenStudy (cggurumanjunath):

(a^2-b^2)=(a+b)(a-b)

OpenStudy (cggurumanjunath):

(((3x^3)^2-(4y^3))^2) a=3x^3 b-4y^3

OpenStudy (cggurumanjunath):

a=3x^3 b=4y^3

OpenStudy (anonymous):

so a - b = -1?

OpenStudy (anonymous):

hello